### v1

# PixCell: A generative foundation model for digital histopathology images

Abstract
The digitization of histology slides has revolutionized pathology, providing massive datasets for cancer diagnosis and research. Contrastive self-supervised and vision-language models have been shown to effectively mine large pathology datasets to learn discriminative representations. On the other hand, generative models, capable of synthesizing realistic and diverse images, present a compelling solution to address unique problems in pathology that involve synthesizing images; overcoming annotated data scarcity, enabling privacy-preserving data sharing, and performing inherently generative tasks, such as virtual staining. We introduce PixCell, the first diffusion-based generative foundation model for histopathology. We train PixCell on PanCan-30M, a vast, diverse dataset derived from 69,184 H&E-stained whole slide images covering various cancer types. We employ a progressive training strategy and a self-supervision-based conditioning that allows us to scale up training without any annotated data. PixCell generates diverse and high-quality images across multiple cancer types, which we find can be used in place of real data to train a self-supervised discriminative model. Synthetic images shared between institutions are subject to fewer regulatory barriers than would be the case with real clinical images. Furthermore, we showcase the ability to precisely control image generation using a small set of annotated images, which can be used for both data augmentation and educational purposes. Testing on a cell segmentation task, a mask-guided PixCell enables targeted data augmentation, improving downstream performance. Finally, we demonstrate PixCell’s ability to use H&E structural staining to infer results from molecular marker studies; we use this capability to infer IHC staining from H&E images. Our trained models are publicly released to accelerate research in computational pathology. {NoHyper}Correspondence at {srikary, agraikos, samaras}@cs.stonybrook.edu
1 Introduction
Generative models, capable of controlled synthesis of realistic tissue images, present a compelling solution to scarcity of rare disease presentations, to annotated data scarcity, to the need for privacy-preserving data sharing, as well as to carry out inherently generative tasks, such as virtual staining. The rise of computational pathology has brought significant advances in automating clinical tasks, such as tumor subtyping and tumor classification. However, developing robust models in this field is often constrained by data-related challenges. For instance, accurate tumor-infiltrating lymphocyte segmentation models, essential in making some clinical predictions, require extensive pixel-level annotations, which are laborious and costly to acquire. Thus, the prevailing approach has been to train powerful self-supervised learning (SSL) models, which benefit from vast amounts of unlabeled data and can be readily adapted to perform downstream tasks.
Self-supervised learning (SSL) models (or vision-language models) are trained to discriminate between different images uni ; virchow1 ; phiconv2 ; hibou (or image-text pairs plip ; quilt ; conch ; musk ), a process that entails learning to extract discriminative representations from a given image. These representations are then utilized to solve a given downstream task, and have been shown to indeed encode useful information for making predictions uni ; phiconv2 ; conch .
Although SSL provides a clear pathway for solving most computational pathology tasks, there are some fundamental drawbacks. Firstly, using a pre-trained self-supervised model to solve any downstream task involves training a prediction ‘head’, which requires a non-trivial amount of high-quality annotated data. Secondly, SSL training benefits from curating diverse datasets to train on Regulatory, data privacy and data ownership issues frequently limit data sharing, therefore, in most cases each institution or consortium ends up training their own SSL model primarily with private data supplemented by a few commonly used public datasets such as the TCGA whole slide imaging dataset weinstein2013cancer . Data sharing challenges limit model performance and can potentially create bias due to the specific demographics of training datasets vaidya2024demographic ; kheiri2025investigation . Finally, tasks where we synthesize a completely new image from a given slide, such as virtual staining (stain translation) boyd2022region ; pati2024accelerating ; li2023adaptive , are difficult to perform with models exclusively trained to discern between different images.
We propose to overcome these limitations using a generative model. We introduce PixCell, the first diffusion-based generative foundation model for histopathology. We employ a Diffusion Transformer peebles2023scalable , and train PixCell on PanCan-30M – an extensive dataset of 30.8 million $1024\times 1024$ px image patches derived from 69,184 H&E-stained whole slide images (WSIs), covering a comprehensive range of cancer types. To effectively scale up training we employ: (a) progressive training, starting from $256\times 256$ patches and gradually increasing the generated image size to $1024\times 1024$ and (b) per-patch guidance, by conditioning PixCell on feature embeddings from a pre-trained self-supervised model (UNI-2h uni ), providing rich image semantics to guide the generation process.
PixCell achieves the state-of-the-art generation quality in $256\times 256$ and $1024\times 1024$ image generation. Benefiting from the pan-cancer training, PixCell significantly outperforms prior models trained on single cancer types cvpr24_ldm ; zoomldm ; inftybrush . To highlight the utility of PixCell-generated data, we train an SSL encoder exclusively on synthetic samples. This encoder achieves comparable performance to a model trained on corresponding real images, which underscores that PixCell’s synthetic data can serve as an effective, privacy-preserving replacement for real data in training downstream models. We release this synthetic dataset and invite the community to experiment with the synthetic images.
PixCell also facilitates fine-grained controllable generation. Using a ControlNet zhang2023adding , we synthesize realistic histopathology images conditioned on cell segmentation masks and demonstrate how these synthetic image-mask pairs can augment the training dataset and improve the performance of a downstream segmentation model. Furthermore, PixCell demonstrates generalization capabilities, successfully synthesizing images with staining techniques unseen during training. We investigate PixCell’s adaptability to stain translation, by fine-tuning the model on a dataset composed of approximately paired H&E and IHC images. PixCell achieves SoTA performance in stain translation, showcasing its potential as a versatile foundation model that can be efficiently adapted to specialized generative tasks using limited additional data.
In summary, our contributions are as follows:
(1) We develop PixCell, the first generative foundation model for histopathology images.
(2) We demonstrate state-of-the-art image generation quality, producing high-fidelity and diverse samples that can be used in place of real images for training self-supervised models.
(3) We explore PixCell’s capability for controllable image generation by conditioning on cell segmentation maps, and showcase its utility through synthetic data augmentation.
(4) We illustrate PixCell’s ability to generalize to different staining techniques, which we use to translate from H&E-stained images to IHC, achieving state-of-the-art results.
(5) We publicly release the PixCell model weights and associated code to foster further research in computational pathology.
2 Method and Data
2.1 Training dataset: PanCan-30M
Our training cohort comprises 69,184 Whole Slide Images (WSIs), aggregated to capture a wide variety of tissue morphologies and cancer subtypes. Inspired by Phikon-v2 phiconv2 , we source the WSIs from major public repositories including The Cancer Genome Atlas (TCGA) weinstein2013cancer , the Clinical Proteomic Tumor Analysis Consortium (CPTAC) edwards2015cptac , the Genotype-Tissue Expression (GTEx) project lonsdale2013genotype , and other public sources chowdhury2023proteogenomic ; national2011reduced , alongside an internal cohort (SBU). The pan-cancer nature of this WSI collection, covering a wide variety of organs, provides a diverse cohort for training a foundation generative model. In Figure 1 we provide a breakdown of the WSIs per organ type used in training.
Figure 1: Distribution of organs in the whole-slide images used for the PanCan-30M dataset.
From these WSIs, we extract 30 million (30,819,977) patches, creating PanCan-30M. Each patch is $1024\times 1024$ pixels in size, and is obtained from slide regions corresponding to 20×20$\times$20 × objective magnification (0.5 microns per pixel). We use the code from DS-MIL li2021dual for both patch extraction and tissue thresholding. The slides included in the dataset are stained with Hematoxylin and Eosin (H&E), and we do not use stain normalization. See Table 1 for an exact breakdown of the data sources for the extracted patches. From this data, we create a separate hold-out test set of 2.3 million image patches, which we call PanCan-Test. In Appendix Figure 15 and Table 9 we provide an exhaustive breakdown of the dataset per organ and data source.
Table 1: Details of the data sources used in the PanCan-30M dataset. We extract the $1024\times 1024$ patches from whole-slide images across various organs from each data source.
Data Source # WSI # Patches
TCGA - diagnostic 11,766 9,192,388
TCGA - fresh frozen 18,310 3,698,529
CPTAC 6,124 2,200,109
SBU 5,698 6,689,518
GTex normal slides 25,713 7,896,145
Other public sources 1,573 1,143,288
Total 69,184 30,819,977
2.2 PixCell diffusion model
Figure 2: Overview of (a) the PixCell-256 model and (b) the PixCell-1024 model. Each model is conditioned on the UNI-2h embeddings that describe the given image.
PixCell is a Diffusion Transformer (DiT) model peebles2023scalable , with its architecture adapted from the PixArt-$\sigma$ framework pixart-sigma . We follow the LDM approach rombach2022high , operating in the latent space of the pre-trained Variational Autoencoder (VAE) from Stable Diffusion 3 sd3 . The VAE first encodes the high-resolution histopathology images from PanCan-30M into compressed latent representations, on which we train the diffusion model. The VAE encoder has a downsampling factor of ×$\times 8$, compressing a $1024\times 1024\times 3$ image to a $128\times 128\times 16$ latent. We opted for the Stable Diffusion 3 VAE, despite the larger latent representation (16-dim in SD3 vs 4-dim in SDXL podellsdxl ), because of the better reconstruction quality in pathology images (average PSNR of 31.79 for SD3 vs 26.93 for SDXL  [^1] PSNR measured on 1000 random $256\times 256$ patches from PanCan-Test.).
To guide PixCell during generation, we condition the model on image embeddings extracted using the self-supervised UNI-2h uni encoder. In natural images, generative foundation models are usually trained on extensive image and text caption pair datasets. For digital pathology images, as there is no large-scale image-caption dataset that we could use to train a generative model on, we can utilize SSL embeddings in place of the text captions cvpr24_ldm . By conditioning on SSL embeddings, we replace the text prompt given in a text-to-image inference with a reference image ’prompt’. The embeddings are integrated into the DiT blocks via cross-attention.
2.2.1 Progressive training
Instead of directly training PixCell on the 30M $1024\times 1024$ image patches, we employ a progressive, multi-stage training strategy with increasing image resolution to optimize efficiency and model performance. This technique is also used for the PixArt family of models pixart-alpha ; pixart-sigma in the natural image domain.
We split training into three stages that utilize the same underlying PanCan-30M dataset of $1024\times 1024$ pixel patches. For the lower-resolution training stages, multiple smaller patches are derived from each original high-resolution patch. To save compute time, we pre-extract VAE features for all $1024\times 1024$ patches in our PanCan-30M dataset, using the pretrained SD-3 VAE sd3 , resulting in $128\times 128\times 16$ dimensional latents. Simultaneously, we extract the $16\times 1536$ dimensional conditioning embedding using the pretrained UNI-2h uni encoder, corresponding to one embedding for each $256\times 256$ patch in the image.
We also create a separate hold-out test set of 2.3 million image patches, which we call PanCan-Test. We utilize the pre-extracted VAE features and UNI-2h embeddings corresponding to the remaining 27.9 million image patches for all training stages of PixCell:
Stage 1 - Low-resolution training (PixCell-256):
We first train a base model on latents corresponding to $256\times 256$ px images. For each $1024\times 1024$ image’s pre-extracted features, we generate 16 non-overlapping $32\times 32\times 16$ crops from its $128\times 128\times 16$ VAE latent. Each of these latent crops is then paired with the corresponding $1\times 1536$ embedding from the original $16\times 1536$ UNI-2h embeddings. This stage is trained for one epoch over the entire dataset, yielding the PixCell-256 model (Figure 2 (a)).
Stage 2 - Intermediate fine-tuning:
The weights from PixCell-256 are used to initialize the model for a second stage of fine-tuning, this time using 512×512 images. From each full-resolution VAE latent, we extract 4 non-overlapping $64\times 64\times 16$ crops. Each $512\times 512$ equivalent latent crop is paired with its corresponding $4\times 1536$ embeddings (representing a $2\times 2$ block of tokens) from the original UNI-2h embeddings extracted. This stage serves to gradually adapt the model before the final high-resolution training.
Stage 3 - High-resolution fine-tuning (PixCell-1024):
Finally, the model from the $512\times 512$ stage is further fine-tuned on latent representations corresponding to full $1024\times 1024$ pixel images for one epoch. The final $1024\times 1024$ training stage directly uses the full pre-extracted $128\times 128\times 16$ VAE features and the complete $16\times 1536$ UNI-2h embeddings. This final stage produces the high-resolution PixCell-1024 model (Figure 2 (b)).
Table 2 summarizes key training parameters, such as the number of iterations and batch sizes, for each stage. Across all stages, we use the AdamW optimizer with a weight decay of $0.03$ and a constant learning rate of $2\times 10^{-5}$. We trained our models on a cluster with several NVIDIA A100 GPUs. More details are given in Appendix A.2.
Table 2: Training details of PixCell. We progressively scale the model training resolution and resources.
Resolution # Images Training steps Batch size # GPUs Conditioning
$256\times 256$ 480 Mil 120,000 4096 16 $1\times 1536$
$512\times 512$ 120 Mil 60,000 1536 24 $4\times 1536$
$1024\times 1024$ 30 Mil 80,000 384 32 $16\times 1536$
2.3 Controllable generation
Figure 3: PixCell-256 with a cell mask ControlNet. The two sources of conditioning allow for fire control over the generated image: the UNI embedding dictates the style of the generated image while the cell mask guides the cell layout.
PixCell generates images conditioned on an embedding from a self-supervised encoder. Within this framework, we can synthesize variations of a given image but lack further control over the synthesized image content. To introduce the ability to controllably generate images, we employ ControlNets zhang2023adding , which add image-level control to pre-trained diffusion models.
We train a ControlNet for PixCell-256 to guide generation with a cell layout mask. To construct a training dataset, we utilize a pre-trained CellViT-SAM-H horst2024cellvit model trained on 0.5 microns per pixel images. We extract cell masks from 10,000 images of all cancer types from PanCan-30M and train the ControlNet with image, UNI embedding, and mask triplets. The diffusion transformer ControlNet copies each layer of the base transformer, adding an intermediate output linear layer that is zero-initialized to combine the base transformer and ControlNet transformer features chen2024pixart . An overview of the ControlNet approach is provided in Figure 3.
2.4 Stain Inference
PixCell is trained to generate H&E-stained histopathology images conditioned on UNI-2h embeddings. However, the UNI encoder has been trained on both H&E and immunohistochemistry (IHC) images. Therefore, we examine whether our model can generalize to produce images with different stains using different conditions.
To test this, we prompt PixCell-256 using an IHC image from the MIST-HER2 dataset li2023adaptive . We extract the UNI-2h embedding of the IHC image and use it to condition the PixCell model. The generated images (Appendix Figure 11) follow the appearance and semantics of the reference IHC patch, showing that, although unseen during training, the model can infer results that would have been obtained with different stains.
IHC stains reveal the spatial distribution of specific molecular markers and are usually significantly more expensive to obtain than H&E. Given that we can generate IHC images, we propose to utilize the pre-trained PixCell model to perform stain translation between H&E and IHC images. Given a dataset of ’roughly’ paired H&E and IHC images, such as MIST li2023adaptive , we introduce a method to generate an IHC-stained image from a given H&E image of the same tissue region. Using a generative model for virtual (re-)staining has the potential to enable further downstream analysis of H&E-stained tissues, without requiring an expensive and likely unavailable IHC staining procedure.
We first extract the UNI embeddings of the paired patches for the H&E and IHC images. We then train a rectified flow model liu2022flow that learns to transform UNI embeddings of H&E images to UNI embeddings of IHC images. We chose a rectified flow instead of a diffusion model because of its direct applicability to learning transformations between two data distributions. Then, for a new H&E image, we first extract the UNI embeddings of each patch, transform them to IHC embeddings, and generate the corresponding IHC image. A main advantage of this conditioning-transformation approach is that we operate entirely on a patch level, which is desirable as most of the paired H&E and IHC datasets lack pixel-perfect alignment.
To further boost the visual quality of the generated IHC data, we show that we can also train a lightweight low-rank adaptation (LoRA) module hu2022lora on top of the PixCell-1024 model, tailored to generating images from the target dataset. Training a LoRA adapter is cheap and only requires the unpaired IHC images. We provide an overview of the proposed stain translation approach in Figure 4.
Figure 4: Overview of our approach for translating H&E stained images to IHC using PixCell-1024. We train an additional H&E-to-IHC model that transforms the UNI conditions from one domain to the other. Adding a low-rank adapter to the PixCell-1024 weights significantly improves the synthesized image quality with minimal additional costs.
3 Experiments
3.1 Image quality evaluation
We first describe the metrics used to evaluate the quality of images generated by PixCell at various resolutions. For $256\times 256$ patches, we compute the following metrics:
Vanilla FID: We measure the Frechet Inception Distance (FID) heusel2017gans between PixCell-generated and real images, using the implementation of Clean-FID parmar2021cleanfid .
Embedding Similarity: We measure the similarity of real and corresponding synthetic images in the embedding space of several pretrained image encoders by calculating the cosine similarity between their respective embeddings. For a robust evaluation, we use UNI-v1 uni , Phicon-v2 phiconv2 , and CONCH-v1 conch as the feature extractors.
For $1024\times 1024$ high-resolution images, we adopt an evaluation strategy consistent with LRDM cvpr24_ldm and MultiDiffusion bar2023multidiffusion . This strategy involves two primary perceptual metrics:
Crop FID: We compare the distribution of $256\times 256$ random crops we extract from the synthesized images against the distribution of real $256\times 256$ patches using the vanilla Inception-v3 network szegedy2016rethinking . CLIP FID: We measure the FID directly between the synthesized images generated by PixCell-1024 and real $1024\times 1024$ images. For this, we extract features using a pretrained CLIP ViT-B/32 model radford2021learning ; KynkaanniemiKAA23 , which is better-suited to evaluating larger images ($>224\times 224$ px).
We sample 10,000 images to measure Vanilla FID and Embedding similarity, following cvpr24_ldm ; pathldm . For large images, we generate 3,000 images for CLIP FID and extract random crops from those images for Crop FID. Unless otherwise specified, we employ the DPM-Solver lu2022dpm with 50 sampling steps and utilize classifier-free guidance ho2022classifier .
3.1.1 Classifier-free guidance scale
To determine the optimal classifier-free guidance scale hyperparameter, denoted as $w$, we perform a grid search across the guidance values $w\in\{1,1.2,1.4,1.6,1.8,2,3,5\}$. For this search, we generate images conditioned on randomly selected UNI-2 embeddings from the training set and measure FID against images from the PanCan-Test set. Table 3 indicates that the best image quality (lowest FID) is obtained for $w=2$ for PixCell-256 and $w=1.2$ for PixCell-1024.
The optimal guidance scale is noticeably lower than the scales regularly used in text-to-image diffusion models for natural images (typically around 6). We attribute this to our choice of conditioning; whereas a single text caption can characterize different images (wider distribution), a UNI embedding corresponds to a much smaller set of images, i.e., the distribution of images the model learns to sample from given a UNI embedding is more concentrated. We use the optimal guidance values for all image generation tasks reported in this paper, unless stated otherwise.
Table 3: Grid search to find the optimal classifier-free guidance scale $w$ based on the FID score.
$w$ 1 1.2 1.4 1.6 1.8 2 3 5
PixCell-256 13.25 12.21 11.16 10.67 10.34 9.65 9.81 10.54
PixCell-1024 8.39 7.92 8.55 9.14 9.72 10.83 15.2 20.59
3.1.2 PixCell-256 results
Figure 5 showcases synthetic samples generated with PixCell-256. We evaluate the image quality of PixCell-256 by comparing its performance against two prior diffusion models: LRDM cvpr24_ldm and ZoomLDM zoomldm . We conduct our comparison by measuring vanilla FID on images from the PanCan-Test set, as well as the TCGA-BRCA dataset, on which both prior works trained their models.
The results, presented in Table 5, show that PixCell-256 outperforms both prior works on the diverse PanCan-Test set, highlighting the robust generalization capabilities and the benefits of PixCell-256’s pan-cancer training approach. On TCGA-BRCA, PixCell lags slightly behind the other methods, which can be attributed to both being trained exclusively on TCGA-BRCA data.
Both previous models (LRDM cvpr24_ldm , ZoomLDM zoomldm ) also use image embeddings from a reference image to generate new images. Therefore, we assess the fidelity and faithfulness to the original image used in generating a patch by measuring the embedding similarity between real and synthetic patches generated from each model on the PanCan-Test set. Table 5 shows that PixCell-256 obtains significantly higher similarity than prior work across multiple pretrained encoders. This further underscores the advantage of our scaled-up pan-cancer training, indicating that PixCell-256 can better preserve essential semantic content from the reference images during synthesis. This ability of PixCell to create new images that retain the properties of a reference image is what we exploit in Section 3.2 to perform self-supervised learning exclusively on synthetic data.
Table 4: FID scores for $256\times 256$ patches generated by different models.
Vanilla FID ↓
Method TCGA-BRCA PanCan-Test
LRDM cvpr24_ldm 6.98 14.16
ZoomLDM zoomldm 6.77 12.04
PixCell-256 8.79 9.65
Table 5: Embedding similarity between real and synthetic $256\times 256$ patches on PanCan-Test.
Embedding similarity ↑
Method UNI-v1 Phicon-v2 CONCH-v1
LRDM cvpr24_ldm 0.408 0.562 0.758
ZoomLDM zoomldm 0.590 0.682 0.825
PixCell-256 0.699 0.827 0.888
Figure 5: Generating variations of a reference image using PixCell-256.
3.1.3 PixCell-1024 results
We evaluate the performance of PixCell-1024 for high-resolution $1024\times 1024$ image generation on the TCGA-BRCA dataset, comparing it against prior methods: LRDM cvpr24_ldm , $\infty$-Brush inftybrush , and ZoomLDM zoomldm . Our assessment considers image quality, using Crop FID and CLIP FID, as well as generation efficiency, measured in generation time per image (seconds). Figure 6 shows synthetic samples generated with PixCell-1024.
As shown in Table 6, PixCell-1024 demonstrates SoTA performance across all evaluated metrics. For image quality, it achieves a Crop FID of 7.92 and a CLIP FID of 0.68, substantially lower than prior work. This signifies that images generated with PixCell-1024 more closely resemble the feature distribution of real images, compared to those from baseline methods. Furthermore, PixCell-1024 achieves the best generation speed, producing a $1024\times 1024$ image in just 2.5 seconds, significantly faster than its counterparts. These findings, coupled with the strong semantic fidelity shown by PixCell-256, highlight the scalability of our PixCell architecture and the effectiveness of the progressive, pan-cancer training strategy from $256\times 256$ up to $1024\times 1024$ resolution.
Figure 6: Generating variations of a reference image using PixCell-1024.
Table 6: Image quality of $1024\times 1024$ patches and generation time on TCGA-BRCA. PixCell-1024 significantly outperforms all previous baselines in sampling time and generation quality.
Method Time / img (s) Crop FID ↓ CLIP FID ↓
LRDM cvpr24_ldm 60 15.51 7.43
$\infty$-Brush inftybrush 30 17.87 3.74
ZoomLDM zoomldm 28 14.94 1.23
PixCell-1024 2.5 7.92 0.68
3.2 Synthetic SSL
In Section 3.1.2, we showed that PixCell-256 synthetic images closely follow the semantics and appearance of the reference image used for extracting the UNI-2h conditioning. This allows us to generate synthetic variants of entire pathology datasets and use them in place of real data. Although this does not guarantee complete data privacy, as generative models have been prone to leaking training samples carlini2023extracting , it is a first step towards enabling synthetic data sharing for model training.
We evaluate the quality of PixCell-generated synthetic images by training an SSL encoder on the synthesized data and comparing its performance against an SSL encoder trained only on real images. To create the real dataset, we sample 1M patches from our internal SBU 12M cohort at $256\times 256$ resolution. We name this dataset Real SBU-1M. Next, using PixCell-256, we generate 1M synthetic images by extracting the UNI embedding and generating a new image for each reference image in Real SBU-1M. We call this dataset Synthetic SBU-1M.
We train a ViT-S DINOv2  dinov2 model for 90 epochs on the Real SBU-1M and Synthetic SBU-1M datasets separately. For evaluation, we use the Thunder benchmark tool thunder to assess the k-NN performance on patch classification. We choose a training-free k-NN evaluation to ensure that the SSL encoder trained on synthetic images is never adapted to real data, making this setting well-suited to measure the effectiveness of the PixCell-generated dataset.
As presented in Table 7, the SSL encoder trained with synthetic images matches the performance of the SSL encoder trained exclusively on real data. The synthetic images can be used as a drop-in replacement of the real images during SSL training without any degradation in downstream performance. We envision that SSL encoders trained on a mix of real and synthetic data will further push the boundaries of self-supervised learning, as they will further scale the training dataset sizes and potentially enable a way to share data between institutions. We release the Synthetic SBU-1M dataset and invite the community to experiment with our synthetic data.
Table 7: k-NN performance of SSL encoders (DINOv2 ViT-S) trained on 1M real and synthetic patches from the SBU dataset. We use the Thunder framework thunder for evaluation on patch classification datasets and report the balanced accuracy.
Patch
Pretraining Dataset BACH bach BreakHis breakhis NCT-CRC nct-crc Camelyon pcam Average
Synthetic SBU-1M 51.6 62.8 90.4 81.5 71.6
Real SBU-1M 50.0 65.1 92.3 79.5 71.7
3.3 Controllable generation
3.3.1 ControlNet training
To generate $256\times 256$ images conditioned on cell segmentation masks, we employed a pre-trained CellViT-SAM-H horst2024cellvit model trained on 20×20$\times$20 × pathology images. We randomly sampled 10,000 images from PanCan-30M and their UNI-2h embeddings, extracted binary cell masks using CellViT, and trained the ControlNet model for 25,000 iterations, using a batch size of 4 and the AdamW optimizer with a learning rate of $1\times 10^{-5}$.
Figure 7: Images generated using the cell mask ControlNet. The synthesized samples follow the appearance of the style image, from which the UNI conditioning is extracted, and the cell layout of the reference mask (columns 1-5). The UNI conditioning from the style image is not sterile from cell information, and when the two significantly contrast, the generated images fail to follow the mask accurately (columns 6-7).
In Figure 7, we showcase images sampled using the trained cell mask ControlNet. The PixCell-256-Cell-ControlNet model allows for finer control over the generated image. By providing the UNI embedding from a reference image and a target cell mask, we have disentangled the appearance (guided by UNI) and cell layout (guided by the mask) of the generated image. However, this disentanglement is not perfect since the UNI-2h conditioning also encodes some information about the cell count and layout. We observe that in some cases, the clashing control signals lead to images that do not exactly follow the given mask (columns 6-7 of Figure 7). Although we can increase the guidance scale of the ControlNet conditioning, this can lead to potentially unwanted artifacts in the generated images or the model completely ignoring the UNI conditioning, as shown in Figure 8.
3.3.2 Targeted data augmentation
Figure 8: Effect of the guidance scale on the cell mask ControlNet generation.
Having trained the cell mask ControlNet, we set up a testbed for data augmentation using synthetic data from our model. We utilize the Lizard dataset graham2021lizard , which provides images at 20×20$\times$20 × magnification alongside their cell segmentation masks. We opt for binary cell segmentation, which our ControlNet is guided by.
As a baseline, we train a segmentation head using features from the last layer of UNI-2h. We use the convolutional segmentation head from the EVA framework kaiko.ai2024eva that combines the extracted features with the input image to better capture cell edges. The segmentation head is trained on the Lizard data splits from GlaS, PanNuke, and CoNSeP, and is evaluated on the CRAG and DigestPath splits. We specifically selected those splits to exacerbate the difference between the training and evaluation data, testing the generalization limits of the extracted features.
We then utilize the PixCell-256-Cell-ControlNet to perform targeted data augmentation on the evaluation data. For each image in the evaluation set, we extract its UNI-2h embedding and pair it with a random training set mask. We then synthesize a new image using the two sources of conditioning. The generated image follows the appearance of the evaluation set images and has a known ground truth cell mask. We use this synthetic data as additional training samples to re-train the segmentation head. Note that this targeted data augmentation scheme does not require ground truth masks from the test set.
For the evaluation, we measure the accuracy, Dice score, and IoU between the predicted and ground truth evaluation masks. We find that with our PixCell augmentations, the accuracy increases from 0.857 →  0.890, Dice score 0.629 →  0.653, and IoU from 0.751 → 0.802. This improvement validates our observation that by training the cell mask ControlNet we attain finer control over the appearance and cell layout of the generated images, which we can then use to synthesize additional training data for downstream tasks.
3.4 Synthesizing IHC stain patterns using H&E images
We use the MIST dataset li2023adaptive to evaluate our stain translation approach. In Appendix Figure 11, we randomly selected patches from the MIST-HER2 slides to extract UNI embeddings from, and show that the base PixCell-256 model can effectively synthesize IHC images without ever being trained explicitly to do so.
To learn the H&E UNI to IHC UNI mapping, we use paired $256\times 256$ patches from MIST to train a rectified flow model liu2022flow . Rectified flows learn to transform samples between two distributions, in this case, H&E UNI embeddings and IHC UNI embeddings. To implement the rectified flow model, we use a simple residual MLP network li2024return and the roughly aligned H&E and IHC patches from the MIST-HER2 dataset. For each pair of images, we extract the two corresponding UNI embeddings and train the rectified flow model to learn the velocity field from H&E to IHC embeddings. During sampling, we start from an H&E UNI embedding and follow the ODE defined by the trained rectified flow to sample a corresponding IHC embedding. The rectified flow model is trained for approximately 50,000 iterations using a batch size of 32 and the AdamW optimizer with learning rate $1\times 10^{-4}$.
In Figure 9 we present the results of this stain translation. We select H&E images from the test set, extract their UNI embeddings, transform them to IHC embeddings with the trained rectified flow model, and generate corresponding IHC images. The resulting samples follow the overall semantics of the ground truth IHC images but noticeably lack in visual quality. Apart from not being trained on IHC data, we attribute the worse quality to the MIST dataset being at a slightly different resolution (0.4661 microns per pixel) than the dataset used for training (0.5 microns per pixel).
To improve the quality of the generated images, we show that we can train a lightweight low-rank adapter hu2022lora for the target dataset. Training a LoRA requires minimal computational resources compared to fine-tuning the entire diffusion model. We train the LoRA on the unpaired MIST-HER2 images for approximately 10,000 iterations, on a single NVIDIA A5000 24GB GPU. The adapter massively improves the generated IHC image quality, without modifying any other component of the stain translation pipeline (Figure 9).
Figure 9: Results for H&E → IHC stain translation using PixCell-1024 with and without LoRA fine-tuning. The quality of the generated images significantly improves with a lightweight low-rank adapter.
Finally, we compare the proposed stain translation to the baseline proposed by the MIST paper, which utilizes a specialized CycleGAN trained for translation on the MIST data. In Table 8 and Figure 10 we provide quantitative and qualitative results, respectively. For quantitative evaluation, we measure the SSIM between the generated and real test images, as well as perceptual metrics (FID, KID, Crop FID). Our LoRA stain translation surpasses the capabilities of the baseline model, which fails to capture critical morphological features in the synthesized IHC images, evident in both the perceptual quantitative metrics and qualitative examples provided.
Our foundation-scale pertaining is the key to effectively performing stain translation. Translating between stains requires a broad understanding of the given tissue features, which we posit is impossible to capture by training only on small datasets. The PixCell model can synthesize immunohistochemistry (IHC) stain patterns from routine H&E slides, reproducing spatial biomarker distributions that normally require separate physical assays for each marker. This virtual staining preserves tumor heterogeneity and microenvironment features within the native tissue context, enabling accurate tumor subtyping and treatment selection. The approach reduces the need for labor-intensive serial sectioning and expensive multiplex assays, rapidly producing pixel-aligned, multiplexed IHC images from a single H&E specimen. Methods of this kind can be used in translational research studies, in low-resource settings, in clinical situations where there is limited tissue availability, or potentially in clinical scenarios as a screening method with molecular confirmation.
Figure 10: Results for H&E →IHC stain translation using PixCell-1024 and CycleGAN baseline. The baseline fails to translate critical features between the H&E and IHC images.
Table 8: Quantitative evaluation on HER2 stain translation.
Method SSIM $\uparrow$ FID $\downarrow$ KID $\downarrow$ Crop FID $\downarrow$
Baseline li2023adaptive 0.1945 54.28 16.0 33.22
PixCell-1024 (no LoRA) 0.1880 67.68 19.1 41.54
PixCell-1024 0.1892 52.32 13.4 20.87
4 Related Work
4.1 Self-Supervised Learning in Computational Pathology
The digitization of histology slides into Whole Slide Images (WSIs) has revolutionized pathology, providing a vast amount of visual data for cancer diagnosis and biomedical research. These gigapixel-scale images capture complex tissue architecture and cellular morphology, offering substantial potential for computational analysis. However, the sheer size of WSIs necessitates processing them at the patch level. A significant challenge arises from the lack of granular, patch-level annotations, making traditional supervised learning approaches difficult. This annotation bottleneck has driven the significant traction of foundation models in digital pathology, particularly self-supervised learning (SSL) at the tile-level uni ; virchow1 ; phiconv2 ; hibou ; kaiko ; gigapath , which allows models to learn meaningful representations from large unlabeled datasets.
SSL enables models to learn generalizable features from the inherent structure of the data itself through pretext tasks. Joint-embedding SSL (JE-SSL) methods, a prominent subclass, aim to learn effective representations by promoting alignment between embeddings of augmented views of the same image while ensuring diversity across the entire dataset. Self-distillation variants within JE-SSL, such as the DINOv2 framework, have demonstrated considerable success in natural image understanding and have been widely adopted in computational pathology (CPath) dino ; ibot ; dinov2 . These tile-based SSL models, including notable examples like UNI uni , Virchow virchow1 , Prov-GigaPath gigapath , and Phikon phiconv2 , are trained on increasingly large and diverse datasets. Early efforts often utilized public repositories like The Cancer Genome Atlas (TCGA), while more recent models leverage extensive proprietary datasets, encompassing millions of slides and billions of tiles. This data often spans multiple institutions, disease states, tissues of origin, and includes a growing variety of staining techniques beyond hematoxylin and eosin (H&E), such as immunohistochemical (IHC) and special stains, as well as data from multiple image magnifications.
Beyond SSL, other discriminative models leverage contrastive learning with paired image-text datasets plip ; quilt ; conch ; musk ; sun2024pathgen . Addressing the gigapixel scale of WSIs, slide-level foundation models have also been developed titan ; gigapath . These operate on larger image regions to capture slide-level context, often by building upon embeddings generated by patch-level encoders (many of which are trained via SSL) in a hierarchical fashion tangle ; gecko . The quality of these patch-level SSL encoders is thus a critical performance gateway for such hierarchical systems.
4.2 Diffusion models
Diffusion models ho2020denoising have established themselves as the main approach for building large-scale generative foundation models. Significant advances in training paradigms improved_ddpm ; song2021scorebased ; lipman2023flow , model architectures ldm ; peebles2023scalable , guidance techniques ho2022classifier ; zhang2023adding ; graikos2023conditional , and sampling strategies ddim ; dpm have enabled these models demonstrate exceptional capabilities in synthesizing high-fidelity, diverse images. Latent Diffusion Models (LDMs) rombach2022high , specifically, reduce the computational cost by compressing the images with a learned encoder-decoder pair pathvae . As also observed by early works on diffusion models dhariwal2021diffusion , accurate conditioning plays a critical role in achieving high-quality image generation.
Examining popular diffusion foundation models in natural images (SD-XL podellsdxl , SD-3 sd3 , PixArt pixart-alpha ; pixart-sigma ), most utilize the LDM approach with a pre-trained encoder and decoder, and train on extensive datasets of billions of image-text pairs to train a text-conditioned diffusion model. Beyond data augmentation TrabuccoDGS24 ; AziziKS0F23 ; gen-sis , the powerful image priors encoded by such foundational generative models have been effectively utilized for a wide range of dense prediction tasks, including image segmentation diffuse_attend and depth estimation Ke_2024_CVPR . A primary factor contributing to the success of diffusion models in natural imaging is the availability of vast, curated collections of image-text pairs. To illustrate, Stable Diffusion 1.4 ldm was trained on LAION-5B schuhmann2022laion , a dataset comprising 5.85 billion such pairs. In comparison, datasets of this scale and nature are not readily available for histopathology.
Despite these pronounced differences, diffusion models have been explored for digital pathology moghadam2023morphology ; muller2023multimodal ; pathldm ; aversa2023diffinfinite . Most are limited to training models on small-scale datasets, partly due to the lack of annotations for training a conditional diffusion model on large datasets. Notably, learned representation-guided diffusion models (LRDMs) cvpr24_ldm provide an alternative by training the generative model conditioned on an embedding produced by a self-supervised encoder. This approach offers a path to scaling up diffusion models in the histopathology domain.
5 Conclusion
We developed PixCell, the first generative foundation model for histopathology images. Trained on PanCan-30M – an extensive dataset derived from 69,184 H&E-stained whole-slide images covering a comprehensive range of cancer types, PixCell demonstrates state-of-the-art image generation quality. It produces high-fidelity and diverse samples that can effectively serve as a replacement for real images in training self-supervised learning models. We explored PixCell’s capability for controllable generation by conditioning on cell segmentation maps and performing synthetic data augmentation for downstream tasks. Finally, we demonstrated PixCell’s adaptability by showcasing its generalization to different staining techniques, which we used to infer IHC staining from H&E images.
We are publicly releasing the PixCell model weights and associated code to foster future research and applications in computational pathology. Our work with PixCell highlights the potential for diffusion-based generative models to serve as true foundational models within the histopathology domain, offering capabilities that extend beyond traditional discriminative approaches. We hope that this contribution will inspire the broader research community to further utilize these models for innovative applications and novel downstream tasks.
Acknowledgments
This research was partially supported by National Institutes of Health (NIH) and National Cancer Institute (NCI) grants 1R21CA258493-01A1, 5U24CA215109, UH3CA225021, U24CA180924, National Science Foundation (NSF) grants IIS-2123920, IIS-2212046, as well as Stony Brook Profund 2022 seed funding and private support from Bob Beals and Betsy Barton. This research used resources of the Argonne Leadership Computing Facility, a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory and is based on research supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program, under Contract No. DE-AC02-06CH11357.
Appendix A Appendix
A.1 Additional results:
In this section, we present additional results from our experiments. In Figure 11, we demonstrate the generalization capabilities of PixCell-256 by generating unseen IHC image patches. In Figure 12, we show example images from our Synthetic-SBU-1M dataset. In Figure 13, we provide additional examples of PixCell-256-Cell-ControlNet generation. In Figure 14, we present additional results of our stain translation approach, and the corresponding result from the CycleGAN baseline of MIST [40].
Figure 11: Using an out-of-distribution IHC image as reference, we generate variations using PixCell-256. Although never trained on IHC data, our model can synthesize images from this unseen domain.
Figure 12: Samples from our Synthetic-SBU-1M dataset. All images in this dataset are generated by providing embeddings from the Real-SBU-1M internal dataset to the PixCell-256 model.
Figure 13: Additional examples of our PixCell-256 cell ControlNet generation.
Figure 14: Additional examples of our PixCell-1024 strain translation and the CycleGAN baseline of MIST [40]
A.2 Training details
We use an NVIDIA DGX A100-based cluster for our training, image quality evaluation, and self-supervised learning experiments. The DGX A100 node comprises eight NVIDIA A100 GPUs providing 320 GB of memory in total. The node contains two AMD EPYC 7742 CPUs with 256 cores and 1 TB of DDR4 memory. The machine uses an SSD for data storage, which offers up to 25 GB/s in bandwidth. In total, we used 3000 node hours for the paper, including exploratory experiments.
For the controllable generation and stain translation experiments, we used NVIDIA RTX A5000 GPUs with 24GB of memory. Training ControlNets and LoRA-based adapters on our PixCell models can be done on as little as a single A5000 GPU, a compute budget trivial in comparison to the PixCell training requirements. We will be publicly releasing the ControlNet and LoRA training scripts.
A.3 Dataset details
In Figure 15, we visualize the distribution of organ types in the PanCan-30M dataset after patch extraction. In Table 9, we present the detailed composition of PanCan-30M, including organ types, data sources, dataset splits, and number of WSIs and patches per data source.
Figure 15: Distribution of organs after patch extraction.
Table 9: Detailed description of the pretraining dataset used in training the PixCell-256 and PixCell-1024 models.
Organ Dataset Split No of WSIs No of patches
Adipose GTEx Adipose - Subcutaneous 978 57,791
GTEx Adipose - Visceral (Omentum) 815 37,763
Adrenal Gland TCGA FFPE acc 227 250,571
TCGA FF acc 96 11,322
GTEx Adrenal Gland 717 263,138
Artery GTEx Artery - Aorta 858 274,691
GTEx Artery - Coronary 662 33,361
GTEx Artery - Tibial 979 61,050
Bladder GTEx Bladder 130 61,382
TCGA FFPE blca 457 459,281
TCGA FF blca 469 79,797
SBU Bladder 57 80,421
Blood-related TCGA FFPE dlbc 44 26,340
TCGA FF dlbc 59 12,427
CPTAC AML 93 164,842
Brain GTEx Brain - Cerebellum 426 154,674
GTEx Brain - Cortex 428 149,441
CPTAC GBM 462 99,722
TCGA FFPE lgg 844 563,253
TCGA FF lgg 728 125,824
TCGA FFPE gbm 860 492,950
TCGA FF gbm 1193 228,187
SBU Brain 15 14,253
Breast GTEx Breast - Mammary Tissue 894 144,924
CPTAC brca 653 162,981
TCGA FFPE brca 1133 735,448
TCGA FF brca 1978 312,557
SBU Breast 1895 1,540,837
Colorectal TCGA FFPE coad 459 267,397
TCGA FF coad 983 196,568
TCGA FFPE read 165 76,709
TCGA FF read 365 72,122
GTEx Colon - Sigmoid 800 281,009
GTEx Colon - Transverse 937 303,477
CPTAC coad 372 126,416
SBU Appendix 199 149,278
SBU Cecum 61 5,544
SBU Colon 354 546,203
SBU Rectum 295 200,305
Gastroesophageal TCGA FFPE esca 158 114,285
TCGA FF esca 238 35,297
TCGA FFPE stad 442 305,349
TCGA FF stad 755 271,865
GTEx Stomach 939 391,985
GTEx Esophagus - Gastroesophageal 787 363,692
GTEx Esophagus - Mucosa 963 211,581
GTEx Esophagus - Muscularis 958 420,428
SBU Esophagus 101 19,980
SBU Gastroesophageal Junction 2 3
SBU Stomach 154 125,812
Organ Dataset Split No of WSIs No of patches
Head and neck CPTAC HNSCC 390 112,431
TCGA FFPE hnsc 472 312,557
TCGA FF hnsc 791 79,785
GTEx Minor Salivary Gland 247 23,017
SBU Epiglottis 1 1
SBU Gingiva 3 3,510
SBU Larynx 171 133,126
SBU Lip 44 49,189
SBU Maxilla 63 46,149
SBU Parotid 53 45,155
SBU Submandibular Gland 9 15,571
SBU Tongue 10 41,621
SBU Tonsil 6 22,393
Heart GTEx Heart - Atrial Appendage 614 246,393
GTEx Heart - Left Ventricle 761 308,624
Kidney GTEx Kidney - Cortex 557 223,111
GTEx Kidney - Medulla 49 19,411
TCGA FFPE kich 121 104,411
TCGA FF kich 205 86,226
TCGA FFPE kirc 519 449,337
TCGA FF kirc 1654 494,982
TCGA FFPE kirp 300 242,237
TCGA FF kirp 473 103,243
CPTAC CCRCC 884 192,805
SBU Kidney 112 489,008
Liver TCGA FFPE chol 39 48,090
TCGA FF chol 71 13,171
TCGA FFPE lihc 379 335,514
TCGA FF lihc 491 82,474
GTEx Liver 610 272,481
SBU Bile Duct 56 1,823
SBU Gallbladder 53 49,131
SBU Liver 259 505,170
Lung TCGA FFPE luad 541 398,649
TCGA FF luad 1067 209,038
TCGA FFPE lusc 512 380,402
TCGA FF lusc 1100 210,390
GTEx Lung 860 297,908
CPTAC LUAD 969 520,881
CPTAC LSCC 1018 488,212
SBU Lung 5 630
others NLST 1225 999,154
Mesothelium TCGA FFPE meso 87 52,567
TCGA FF meso 88 14,085
SBU Peritoneum 1 1
Muscle - skeletal CPTAC SAR 305 95,257
TCGA FFPE sarc 600 639,211
TCGA FF sarc 290 48,382
GTEx Muscle - Skeletal 1001 356,578
SBU Stomach 154 125,812
Organ Dataset Split No of WSIs No of patches
Other Types TCGA FFPE pcpg 196 184,298
TCGA FF pcpg 189 25,688
GTEx Nerve - Tibial 975 95,623
SBU Bone 29 51,297
SBU Cervix 2 11,402
SBU Fallopian Tube 7 320
SBU Lymph Node 238 442,527
SBU Mediastinum 18 22,158
SBU Mesentery 9 1
SBU Negative 11 2,303
SBU Omentum 10 148
SBU Peripheral Nerve 6 1,076
SBU Soft Tissue 61 143,760
SBU Spinal Cord 5 1,387
SBU Unknown 102 106,210
SBU Ureter 17 27,609
Ovary TCGA FFPE ov 107 102,005
TCGA FF ov 1374 372,215
GTEx Ovary 255 105,532
SBU Ovary 17 15,135
others PTRC-HGSOC 348 144,134
Pancreas TCGA FFPE paad 209 168,177
TCGA FF paad 257 42,614
GTEx Pancreas 865 374,094
CPTAC PDA 567 98,216
SBU Ampulla 142 90,211
SBU Pancreas 449 473,837
Pituitary Gland GTEx Pituitary 428 81,116
Prostate GTEx Prostate 599 262,357
TCGA FFPE prad 449 363,268
TCGA FF prad 723 139,815
SBU Prostate 5 18,765
Skin TCGA FFPE uvm 80 46,105
TCGA FF uvm 70 8,454
TCGA FFPE skcm 475 387,694
TCGA FF skcm 475 75,806
GTEx Skin - Not Sun Exposed 818 333,848
GTEx Skin - Sun Exposed 1001 359,139
CPTAC CM 411 138,346
SBU Skin 116 122,049
Small intestine GTEx Small Intestine - Terminal 798 180,567
SBU Duodenum 56 22,758
SBU ileum 12 9,905
SBU Small Bowel 99 129,820
Spleen GTEx Spleen 874 376,920
SBU Spleen 5 24,999
Testis TCGA FFPE tgct 254 240,765
TCGA FF tgct 159 27,182
GTEx Testis 592 208,431
SBU Stomach 154 125,812
Organ Dataset Split No of WSIs No of patches
Thyroid TCGA FFPE thca 519 398,380
TCGA FF thca 639 94,571
GTEx Thyroid 902 283,835
SBU Thyroid 285 796,642
Uterus GTEx Cervix - Ectocervix 38 13,062
GTEx Cervix - Endocervix 42 14,578
GTEx Fallopian Tube 43 4,979
GTEx Uterus 237 106,294
GTEx Vagina 276 137,860
TCGA FFPE ucs 91 99,858
TCGA FF ucs 63 9,524
TCGA FFPE cesc 279 175,492
TCGA FF cesc 325 49,798
TCGA FFPE ucec 566 601,015
TCGA FF ucec 805 143,477
SBU Uterus 16 82,294
SBU Vagina 1 1